# Local Map Construction with SDMap: A Comprehensive Survey

### v2

@@ Abstract
1 Introduction
In intelligent driving, the construction of a local map is crucial for providing real-time awareness of the vehicle’s surroundings, enabling precise navigation. It reflects the current state of the environment using sensor data and helps in predicting the movement of surrounding objects. By offering real-time updates, local maps support dynamic decision-making and control, even in unmapped areas. They enhance robustness, reduce computational load, and ensure that the vehicle can safely adapt to rapidly changing conditions. Local maps provide accurate static information such as road geometries and traffic sign locations, serving as prior knowledge that enhances the robustness of perception systems in complex environments. Local map perception systems are expected to achieve high-precision detection in all weather and lighting conditions, as any instance of missed or incorrect detection can result in significant consequences.
@@ 1 Introduction
Following this introduction, the article unfolds as follows: Section II 2 provides a comprehensive review of the pertinent literature, elucidating the general pipeline for local map construction utilizing SDMap. Section III 3 delves into the concept of SDMap, its advantages, and its application scenarios within the task of local map construction. Additionally, we present an overview of the public datasets commonly employed and the evaluation metrics pertinent to tasks of local map perception. Section IV 4 scrutinizes the representation, encoding, and fusion methods of multi-source information, which are crucial for the generation of real-time local maps. This paper further discusses the underlying principles, architecture, and performance of these methods. Section V 5 identifies and discusses key challenges and potential research directions associated with leveraging SDMap as prior information for local map construction. Finally, Section VI 6 synthesizes our findings and offers perspectives on future research endeavors.
@@ 2.2 Standard Definition Map
2.3 Datasets
In the field of BEV local map construction, commonly used datasets include KITTI[9], nuScenes[10], ApolloScape[11], Argoverse[12], Openlane[13], and Waymo[14] Open Dataset (Table 1).
Table 1: Datasets for BEV local map construction
@@ 2.3 Datasets
nuScenes[10], released by Motional, includes data from six cameras, five radars, one lidar, IMU, and GPS, suitable for urban traffic scenarios under various weather and lighting conditions. The nuScenes dataset[10] features comprehensive annotations for intelligent driving, including 3D bounding boxes for 23 classes of objects and 8 attributes. Each scene, lasting 20 seconds, is fully annotated with detailed 3D bounding boxes. The annotation process involves sampling keyframes at 2Hz and is performed by expert annotators to ensure accuracy. The dataset also provides continuous annotations for objects throughout each scene if they are covered by lidar or radar points.
ApolloScape[11], released by Baidu, offers high-precision 3D annotated data covering various urban road scenes, suitable for tasks like lane detection and semantic segmentation. The ApolloScape dataset[11] includes per-pixel semantic segmentation annotations with a unique scale and complexity. It features over 140K images with pixel-level semantic masks and is planned to scale up to 1 million images. The dataset is annotated using an interactive and efficient labeling pipeline that integrates 2D and 3D labeling stages, leveraging high-quality 3D point clouds to enhance the annotation process. Each image is tagged with high-accuracy pose information, and the dataset contains fine-grain labeling of lane markings, making it the first open dataset with 3D annotations for street views.
@@
2.4 Common Evaluation Metrics
@@
2.4.1 Metrics for Lane Extraction
@@
2.4.2 Metrics for Topology Reasoning
@@ 3.3.2 Representation of Vectors
3.3.3 Encoding of Other Information
@@
4 Multimodal Fusion Method
@@
Additionally, several monocular 3D lane detection methods, such as[41], [42], and[43], focus solely on visual images as input. Numerous models, including[6], [5], [44], [45], [46], [47], [48], [49], [50], [51], [52], [53], [54], [49], [55] and [56], also rely solely on visual images. On the other hand, HDMapNet[57], a representative multimodal method, integrates point clouds by encoding these features and predicting vectorized map elements in BEV, achieving effective fusion of multi-sensor data. Furthermore, other models, such as[58], [59], [60], [61], [62], [63], and [64], incorporate point cloud data as additional input. Fig. 4 illustrates the development trends in local map construction over recent years. Considering the cost of constructing high-precision maps, Maplite 2.0[33] was the first to introduce SDMap into local map perception tasks. SMERF[2] and P-MapNet[31] combine the feature representation of SDMap with camera input features using a multi-head cross-attention mechanism, enabling more effective lane topology inference. To achieve an effective fusion of visual BEV features and SDMap semantic information, BLOS-BEV[65] has explored various feature fusion methods. Additionally, methods such as PriorLane[35], FlexMap[66], Bayesian[67], TopoLogic[68], LGMap[69], MapVision[36], LSM (Leveraging SD Map to Assist the OpenLane Topology)[70], UniHDMap[71], RoadPainter[72], and EORN[73] integrate SDMap priors to support local map construction, a trend that is gradually gaining traction (Table 2).
Table 2: Performance of using SDmap method on dataset OpenLane-V2
@@
PriorLane[35] primarily consists of three successive components: the MiT block, the KEA module, and the FT block. According to the structure depicted in Fig. 6, Unlike typical ViTs[25], the MiT block can generate hierarchical features due to its design of overlapping patch merging and efficient self-attention mechanisms. The Knowledge Embedding Alignment (KEA) module is essential for spatially aligning the knowledge embeddings with the coarse vehicle positions in the image features. The knowledge encoder layer and the fusion encoder layer work together to form the fusion transformer. The FT block introduces a novel approach to detection transformer pre-training by leveraging implicit vision prompts, which in PriorLane[35] are generated from explicit prior knowledge, and notably, no annotated labels are used for prompt training.
@@
MapLite 2.0[33] converts a coarse road graph into an HDmap prior and a rasterized BEV representation. The system generates additional BEV images from sensor inputs, which are processed using the CNN trained on existing HDMap data to predict semantic labels. These semantic segmentation results are then converted into distance transforms specific to each label. A structured estimator simultaneously updates the local map and integrates the SDMap prior. This approach enables the online generation of HDMap, relying solely on SDMap and onboard sensor data. As a result, autonomous vehicles can navigate unmapped areas, while the system also detects environmental changes that could signal the need for map updates.
@@
To create BEV features, MapLite[34] demonstrates that point clouds from lidar sensors are gathered over time and paired with camera images (Fig. 7). These points are then projected into a BEV frame by dividing the space around the vehicle into a grid that captures intensity, height, and RGB values for each grid cell. A ResNet-101-based semantic segmentation model is utilized to predict the positions of important map objects such as lane markings and road boundaries. This model is fine-tuned via transfer learning from a DeepLabv3 network[76] pre-trained on the COCO dataset[77]. After segmentation, the output is converted into signed distance transforms, which handle occlusions and remove artifacts caused by missing data.
@@
Many existing algorithms construct BEV features using multi-perspective images and employ multi-task heads to delineate road centerlines, boundary lines, pedestrian crossings, and other areas. However, performance tends to decline at the far ends of roads, and challenges arise when the primary subject in the image is occluded. To tackle these issues, It can be seen from Fig. 11 that MapVision[36] incorporates both multi-perspective images and SDMap as inputs. SDMap is utilized as a supplementary element, enhancing the understanding of road topology from a BEV perspective while providing map priors over extended distances.
@@
RoadPainter[72] processes surrounding images and optionally incorporates SDMap to generate BEV features. These features are used to detect lane centerlines as geometric points in the BEV space and determine their associations, improving the accuracy and reliability of lane topology. To enhance this process, RoadPainter[72] utilizes an SDMap interaction module that augments BEV features with beyond-visual-range data and road shape priors extracted from the SDMap. The SDMap is vectorized and converted into BEV features by filling each grid cell with the corresponding semantic type embedding if occupied.
For extracting lane centerlines, RoadPainter[72] employs a transformer decoder with a hybrid attention layer, which integrates masked cross-attention, deformable cross-attention, and self-attention. Additionally, it uses a topology head based on centerline instance queries and positional embeddings to achieve accurate topological associations.
@@
To refine the detected centerline points, RoadPainter[72] introduces a points-mask optimization module as illustrated in Fig. 17, consisting of two submodules: points-guided mask generation and points-mask fusion. This method addresses accuracy issues, particularly in areas of high curvature, by leveraging detected centerline points to guide the generation of mask queries. The first stage, mask points sampling, selects a set of points from the generated mask. In the second stage, points fusion combines these sampled mask points with the detected centerline points to produce refined centerline positions. This refinement approach ensures more precise and reliable lane topology reasoning.
EORN[73] centers on utilizing lightweight and scalable SDMap to develop online vectorized HDmap representations. Initially, the integration of rasterized SDMap into various online mapping architectures is explored. A key takeaway from this process is that SDMap encoders are model-agnostic, meaning they can be quickly adapted to new architectures that use BEV encoders. Fig. 18 demonstrates that incorporating SDMap as priors in online mapping tasks significantly accelerates convergence and improves the performance of tasks like online centerline perception.
@@ 4.2 Fusion Methods
5 Challenges and Future Prospects

### v3

@@ Abstract
Index Terms:
Intelligent Driving, local map construction, standard definition map, multimodal fusion.
I Introduction
In intelligent driving system, the construction of a local map is crucial for providing real-time awareness of the vehicle’s surroundings, enabling precise navigation. It reflects the current state of the environment using sensor data and helps in predicting the movement of surrounding objects. By offering real-time updates, local maps support dynamic decision-making and control, even in unmapped areas. They enhance robustness, reduce computational load, and ensure that the vehicle can safely adapt to rapidly changing conditions. Local maps provide accurate static information such as road geometries and traffic sign locations, serving as prior knowledge that enhances the robustness of perception systems in complex environments. Local map perception systems are expected to achieve high-precision detection in all weather and lighting conditions, as any instance of missed or incorrect detection can result in significant consequences.
@@ I Introduction
Following this introduction, the article unfolds as follows: Section II provides a comprehensive review of the pertinent literature, elucidating the general pipeline for local map construction utilizing SDMap. Section III delves into the concept of SDMap, its advantages, and its application scenarios within the task of local map construction. Additionally, we present an overview of the public datasets commonly employed and the evaluation metrics pertinent to tasks of local map perception. Section IV scrutinizes the representation, encoding, and fusion methods of multi-source information, which are crucial for the generation of real-time local maps. This paper further discusses the underlying principles, architecture, and performance of these methods. Section V identifies and discusses key challenges and potential research directions associated with leveraging SDMap as prior information for local map construction. Finally, Section VI synthesizes our findings and offers perspectives on future research endeavors.
@@ II-B Standard Definition Map
II-C Datasets
In the field of BEV local map construction, commonly used datasets include KITTI[9], nuScenes[10], ApolloScape[11], Argoverse[12], Openlane[13], and Waymo[14] Open Dataset (Table I).
TABLE I: Datasets for BEV local map construction
@@ II-C Datasets
nuScenes[10], released by Motional, includes data from six cameras, five radars, one lidar, IMU, and GPS, suitable for urban traffic scenarios under various weather and lighting conditions. The nuScenes dataset[10] features comprehensive annotations for intelligent driving, including 3D bounding boxes for 23 classes of objects and 8 attributes. Each scene, lasting 20 seconds, is fully annotated with detailed 3D bounding boxes. The annotation process involves sampling keyframes at 2Hz and is performed by expert annotators to ensure accuracy. The dataset additionally includes continuous annotations for objects across each scene, provided they are captured by lidar or radar points.
ApolloScape[11], released by Baidu, offers high-precision 3D annotated data covering various urban road scenes, suitable for tasks like lane detection and semantic segmentation. The ApolloScape dataset[11] includes per-pixel semantic segmentation annotations with a unique scale and complexity. It features over 140K images with pixel-level semantic masks and is planned to scale up to 1 million images. The dataset is annotated using an interactive and efficient labeling pipeline that integrates 2D and 3D labeling stages, leveraging high-quality 3D point clouds to enhance the annotation process. Each image is accompanied by highly accurate pose information, and the dataset features detailed labeling of lane markings, making it the first publicly available dataset to offer 3D annotations for street views.
@@
II-D Common Evaluation Metrics
@@
II-D1 Metrics for Lane Extraction
@@
II-D2 Metrics for Topology Reasoning
@@ III-C2 Representation of Vectors
III-C3 Encoding of Other Information
@@
IV Multimodal Fusion Method
@@
Additionally, several monocular 3D lane detection methods, such as[41], [42], and[43], focus solely on visual images as input. Numerous models, including[6], [5], [44], [45], [46], [47], [48], [49], [50], [51], [52], [53], [54], [49], [55] and [56], also rely solely on visual images. On the other hand, HDMapNet[57], a representative multimodal method, integrates point clouds by encoding these features and predicting vectorized map elements in BEV, achieving effective fusion of multi-sensor data. Furthermore, other models, such as[58], [59], [60], [61], [62], [63], and [64], incorporate point cloud data as additional input. Fig. 4 illustrates the development trends in local map construction over recent years. Considering the cost of constructing high-precision maps, Maplite 2.0[33] was the first to introduce SDMap into local map perception tasks. SMERF[2] and P-MapNet[31] combine the feature representation of SDMap with camera input features using a multi-head cross-attention mechanism, enabling more effective lane topology inference. To achieve an effective fusion of visual BEV features and SDMap semantic information, BLOS-BEV[65] has explored various feature fusion methods. Additionally, methods such as PriorLane[35], FlexMap[66], Bayesian[67], TopoLogic[68], LGMap[69], MapVision[36], LSM (Leveraging SD Map to Assist the OpenLane Topology)[70], UniHDMap[71], RoadPainter[72], and EORN[73] integrate SDMap priors to support local map construction, a trend that is gradually gaining traction (Table II).
TABLE II: Performance of using SDmap method on dataset OpenLane-V2
@@
PriorLane[35] primarily comprises three sequential components: the MiT block, the KEA module, and the FT block. According to the structure depicted in Fig. 6, Unlike typical ViTs[25], the MiT block can generate hierarchical features due to its design of overlapping patch merging and efficient self-attention mechanisms. The Knowledge Embedding Alignment (KEA) module is essential for spatially aligning the knowledge embeddings with the coarse vehicle positions in the image features. The fusion transformer is composed of the knowledge encoder layer and the fusion encoder layer working in tandem. The FT block presents an innovative method for detection transformer pre-training, utilizing implicit vision prompts, which in PriorLane[35] are derived from explicit prior knowledge, with the notable feature that no annotated labels are required for prompt training.
@@
MapLite 2.0[33] converts a coarse road graph into an HDmap prior and a rasterized BEV representation. Additionally, the system produces BEV images derived from sensor inputs, which are processed using the CNN trained on existing HDMap data to predict semantic labels. These semantic segmentation results are then converted into distance transforms specific to each label. A structured estimator simultaneously updates the local map and integrates the SDMap prior. This approach enables the online generation of HDMap, relying solely on SDMap and onboard sensor data. As a result, autonomous vehicles can navigate unmapped areas, while the system also detects environmental changes that could signal the need for map updates.
@@
To create BEV features, MapLite[34] demonstrates that point clouds from lidar sensors are gathered over time and paired with camera images (Fig. 7). The points are subsequently projected into a BEV frame by partitioning the space surrounding the vehicle into a grid, which captures intensity information, height, and RGB values for each grid cell. A ResNet-101-based semantic segmentation model is utilized to predict the positions of important map objects such as lane markings and road boundaries. This model is fine-tuned via transfer learning from a DeepLabv3 network[76] pre-trained on the COCO dataset[77]. After segmentation, the output is converted into signed distance transforms, which handle occlusions and remove artifacts caused by missing data.
@@
Many current algorithms build BEV features from multi-perspective images and utilize multi-task heads to identify road centerlines, boundary lines, and other areas. However, performance often diminishes at the far ends of roads and faces difficulties when the main subject in the image is obscured. To tackle these issues, It can be seen from Fig. 11 that MapVision[36] incorporates both multi-perspective images and SDMap as inputs. SDMap is utilized as a supplementary element, enhancing the understanding of road topology from a BEV perspective while providing map priors over extended distances.
@@
RoadPainter[72] processes surrounding images and optionally incorporates SDMap to generate BEV features. These features are used to detect lane centerlines as geometric points in the BEV space and determine their associations, improving the accuracy and reliability of lane topology. RoadPainter[72] enhances this process by utilizing an SDMap interaction module, which augments BEV features with road shape priors and beyond-visual-range data extracted from the SDMap. The SDMap is vectorized and converted into BEV features by filling each grid cell with the corresponding semantic type embedding if occupied.
For extracting lane centerlines, RoadPainter[72] employs a transformer decoder with a hybrid attention layer, which integrates masked cross-attention, deformable cross-attention, and self-attention. To achieve accurate topological associations, it additionally employs a topology head that utilizes centerline instance queries and positional embeddings.
@@
To refine the detected centerline points, RoadPainter[72] introduces a points-mask optimization module as illustrated in Fig. 17, consisting of two submodules: points-guided mask generation and points-mask fusion. By leveraging detected centerline points to guide the generation of mask queries, this method addresses accuracy issues, particularly in areas of high curvature. The first stage, mask points sampling, selects a set of points from the generated mask. In the second stage, points fusion combines these sampled mask points with the detected centerline points to produce refined centerline positions. This refinement approach ensures more precise and reliable lane topology reasoning.
EORN[73] centers on utilizing lightweight and scalable SDMap to develop online vectorized HDmap representations. Initially, the integration of rasterized SDMap into various online mapping architectures is explored. A key takeaway from this process is that SDMap encoders are model-agnostic, allowing for rapid adaptation to new architectures that use BEV encoders. Fig. 18 demonstrates that incorporating SDMap as priors in online mapping tasks significantly accelerates convergence and improves the performance of tasks like online centerline perception.
@@ IV-B Fusion Methods
V Challenges and Future Prospects